The area of a square garedn is 242 m^2. How long is the diagonal?
Found 2 solutions by mananth, nyc_function:
Answer by mananth(16946)

area = 242 sq mtr
side = sqrt 242
diagonal^2 = side1^2 +side2^2
diagonal^2 = 2side^2 since it is a square
diagonal^2= 2 (sqrt242)^2
diagonal^2= 2*242
= 484
diagonal = sqrt 484 = 22







Answer by nyc_function(2741)

Let s = side of square.
Area = side^2
242 = side^2
sqrt{242} = side
We now use the Pythagorean Theorem.
[sqrt{242}]^2 + [sqrt{242}]^2 = d^2
Let d = diagonal.
242 + 242 = d^2
484 = d^2
sqrt{484} = sqrt{d^2}
22 = d
Your diagonal is 22.